# Exponential quantum advantages in learning quantum observables from classical data

### v1

@@ A.1 Classical hardness
The main idea of the proof is to show that if there exists a classical algorithm which learns by Def. 1 the time evolution problem for any local Hamiltonian, then such algorithm could decide any BQP language on average implying $\mathrm{BQP}\subseteq\mathrm{HeurP}^{\mathcal{D}}/\mathrm{poly}$, for any $\mathcal{D}$. As in Lemma 3 in [GD23] it is proved that, under widely believed assumptions, for any BQP complete language $\mathcal{L}$ there always exists a distribution $\mathcal{D}$ such that $(\mathcal{L},\mathcal{D})\not\subset\mathrm{HeurP}/\mathrm{poly}$, such classical algorithm cannot exist. Be $\{U^{n}_{BQP}\}_{n}$ a family of quantum circuit which decides the BQP complete language $\mathcal{L}$, one circuit per size. Precisely, that means that for any $x\in\{0,1\}^{n}$ measuring the first qubit in the computational basis on the state $U^{n}_{BQP}\ket{\bm{x}}$ will output 1 (or 0) with probability greater that 2/3 if $\bm{x}\in\mathcal{L}$ ($\bm{x}\not\in\mathcal{L}$). It is obvious that the the quantum model $f^{O^{\prime}}=\text{Tr}[\rho(\bm{x})O^{\prime}]$ correctly decides every $\bm{x}\in\mathcal{L}$ if $\rho(\bm{x})=U^{n}_{BQP}\ket{\bm{x}}\bra{\bm{x}}(U^{n}_{BQP})^{\dagger}$ and $O^{\prime}=Z\otimes\underbrace{I\otimes...\otimes I}_{n-1}$. In fact measuring the first qubit of $U^{n}_{BQP}\ket{\bm{x}}$:
@@
Therefore, as $f^{O^{\prime}}(\bm{x})>0$ if $\bm{x}\in\mathcal{L}$ and $f^{O^{\prime}}(\bm{x})<0$ if $\bm{x}\notin\mathcal{L}$ such quantum model could efficiently decide the language $\mathcal{L}$. By Lemma 3 for every $n$ there exist a 4-local Hamiltonian $H_{\mathrm{\mathrm{hard}}}$ such that evolving the state |$\ket{\bm{x}}$ for a time $t=\pi$ under it will produce the state $U^{n}_{\mathrm{BQP}}\ket{\bm{x}}$ (on the work register). Since such constructed $H_{\mathrm{hard}}$ is local, its time evolution can be efficiently implemented on a quantum computer. Working towards a contradiction, assume now that there exists an algorithm $A_{c}$ which learns $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ with respect to the distribution $\mathcal{D}$ by Def.1. Then it means that there exists a randomized classical algorithm which makes use of data such that $h_{cl}(.)=A_{c}(H_{\mathrm{hard}},\mathcal{T}^{\bm{\alpha}},.)$ satisfies $\forall\epsilon\geq 0$ and $\forall f^{\bm{\alpha}}\in\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ the condition:
$\mathbb{E}_{\bm{x}\sim\mathcal{D}}\big{[}\;|f^{{\bm{\alpha}}}(\bm{x})-h_{cl}(\bm{x})|^{2}\;\big{]}\leq\epsilon$ (22)
with high probability over the choice of the training dataset and the internal randomization of $A_{c}$. We want now to show that the classical model $h_{cl}$ would be able to decide the language $\mathcal{L}$ on average with respect to $\mathcal{D}$. Firstly notice that as the algorithm $A_{c}$ succeeds with high probability, there exists a fixed dataset $\mathcal{T}^{\prime\bm{\alpha}}$ and a fixed bitrsting $r^{\prime}$ for its internal randomization such that $h^{\prime}_{cl}(.)=A_{c}(H_{\mathrm{hard}},\mathcal{T}^{\prime\bm{\alpha}},r^{\prime},.)$ satisfies Eq.(22). Throughout the reminder of the proof let us fix such dataset $\mathcal{T}^{\prime\bm{\alpha}}$ and random bitstring $r^{\prime}$. Also, since Def. 1 assumes Eq. (22) is satisfied for every concept ${\bm{\alpha}}$, such an $A_{c}$ would be able to evaluate on avarage the particular concept ${\bm{\alpha}}^{\prime}$ which corresponds to the observable $O({\bm{\alpha}}^{\prime})=O^{\prime}=Z\otimes I\otimes...\otimes I$. We now show that such classical algorithm would be able to decide the language $\mathcal{L}$ in $\mathrm{HeurP}/\text{poly}$ with respect to the distribution $\mathcal{D}$. Remember that a language $L$ and distribution $D$ is in $\mathrm{HeurP}/\text{poly}$ if there exist a classical algorithm $\mathcal{A}$ and a polynomial-size advice string $\omega$ such that
$\text{Pr}_{\bm{x}\sim\mathcal{D}}[\mathcal{A}(\omega,\bm{x})\neq L(\bm{x})]\leq\epsilon$ (23)
where $L(\bm{x})=1$ if $\bm{x}\in L$ and 0 otherwise. Notice that Eq.(23) is equivalent to
$\sum_{\bm{x}\in\mathcal{X}_{err}}p(\bm{x})\leq\epsilon$ (24)
where $\mathcal{X}_{err}$ is the set of input $\bm{x}$ such that $[\mathcal{A}(\omega,\bm{x})\neq L(\bm{x})]$, i.e. $\mathcal{X}_{err}=\{\bm{x}\in\{0,1\}^{n}\;\;|\;\;[\mathcal{A}(\omega,x)\neq L(\bm{x})\}$, and $p(\bm{x})$ is the probability associated to the input $\bm{x}$ with respect to the distribution $\mathcal{D}$.
[…]
$\epsilon^{\prime}$ $\geq\mathbb{E}_{\bm{x}\sim\mathcal{D}}\big{[}\;|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}\;\big{]}$ (25)
$\geq\sum_{\bm{x}}|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}p(\bm{x})$ (26)
$\geq\sum_{\bm{x}\in\mathcal{X}}|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}p(\bm{x})+\sum_{\bm{x}\not\in\mathcal{X}}|f^{{\bm{\alpha}}}(\bm{x})-h^{\prime}_{cl}(\bm{x})|^{2}p(\bm{x})$ (27)
$\geq 0.09\sum_{x\in\mathcal{X}}p(\bm{x})$ (28)
From last equation we obtain that $\sum_{\bm{x}\in\mathcal{X}}p(\bm{x})\leq\epsilon$. Notice that the classical algorithm $h^{\prime}_{cl}$ is such that $h^{\prime}_{cl}(\bm{x})>0$ if $x\in\mathcal{L}$ and $h^{\prime}_{cl}(\bm{x})<0$ if $\bm{x}\in\mathcal{L}$ for every $\bm{x}\in\{0,1\}^{n}/\mathcal{X}$. Then, the condition $\sum_{\bm{x}\in\mathcal{X}}p(\bm{x})\leq\epsilon$ is exactly the defining condition of $\mathrm{HeurP}/\mathrm{poly}$ in Eq.(24). This implies that the classical algorithm $A_{cl}$ using as advice $\mathcal{T}^{\prime\bm{\alpha}}$ and $r^{\prime}$ is able to decide the language $\mathcal{L}$ in $\mathrm{HeurP}/\mathrm{poly}$ with respect to the distribution $\mathcal{D}$. This concludes the proof. ∎
A.2 Quantum Learnability
In this Section we prove that Algorithm 1 satisfies the learning condition in Def. 1 using only polynomial resources in both time and samples. As a first step, notice that the quanutm states $\rho_{H_{\mathrm{hard}}}(\bm{x})$ in Eq. (1) of $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ are easily preparable on a quantum computer.
@@ A.2 Quantum Learnability
Since the quantum algorithm can efficiently prepare time evolved states of local Hamiltonians, the quantum states $\rho_{H_{\mathrm{hard}}}(\bm{x})=U\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}$ with $U=e^{iH_{\mathrm{\mathrm{hard}}}\pi}$ can efficiently be prepared on a quantum computer. In order to demonstrate the quantum learnability of $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$, all that remains is to rigorously bound the sample and time complexity of Algorithm 1 in the following Theorem 6. As the bound obtained on the sample complexity derives from the generalization bound of the LASSO algorithm [MRT18, LHT+24], we first repeat this result.
Theorem 5 (Theorem 11.16 in [MRT18]).
@@
Let $\mathcal{X}\subseteq\mathbb{R}^{A}$ and $\mathcal{C}=\{\bm{x}\in\mathcal{X}\mapsto\vec{w}\cdot\bm{x}:\|\vec{w}\|_{1}\leq B\}$. Let $\mathcal{S}=((\bm{x}_{1},y_{1}),\dots,(\bm{x}_{N},y_{N}))\in(\mathcal{X}\times\mathcal{Y})^{N}$. Let $\mathcal{D}$ denote a distribution over $\mathcal{X}\times\mathcal{Y}$ according to which the training data $\mathcal{S}$ is drawn. Assume that there exists $r_{\infty}>0$ such that for all $\bm{x}\in\mathcal{X}$, $\|x\|_{\infty}\leq r_{\infty}$ and $M>0$ such that $|h(\bm{x})-y|\leq M$ for all $(\bm{x},y)\in\mathcal{X}\times\mathcal{Y}$. Then, for any $\delta>0$, with probability at least $1-\delta$, each of the following inequalities holds for all $h\in\mathcal{C}$:
$\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}[|h(\bm{x})-y|^{2}]:=R(h)\leq\hat{R}_{\mathcal{S}}(h)+2r_{\infty}BM\sqrt{2\log(2|A|)N}+M\sqrt{2\log(1/\delta)}/{2N}$
@@
We are now ready to state Theorem 6 which provides a rigorous guarantee on the number of samples and the time complexity required by Algorithm 1.
@@
Given $n$, $\delta,\epsilon\geq 0$ and a training data set $\mathcal{T}^{\bm{\alpha}}_{N}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ of size
$N=\mathcal{O}\left(\frac{\log(n/\delta)n}{\epsilon}\right)$ (29)
where $x_{\ell}$ is sampled from an unknown distribution $\mathcal{D}$ over $x\in\{0,1\}^{n}$ and $|y_{\ell}-\text{Tr}(O({\bm{\alpha}})\rho(\bm{x}_{\ell}))|\leq\epsilon$ for any geometrically local observable $O({\bm{\alpha}})=\sum_{i}^{m}{\bm{\alpha}}_{i}P_{i}$ with $P_{i}\in\{I,X,Y,Z\}^{\otimes n}$ and ${\bm{\alpha}}\in[-1,1]^{m}$, and $\rho(\bm{x})=U(\bm{x})\ket{0}\bra{0}U^{\dagger}(\bm{x})$. Then there exists a quantum algorithm $\mathcal{A}_{q}(\mathcal{T}_{N},x)=h(\bm{x})$ such that:
@@
with probability at least $1-\delta$. The computational time of the quantum algorithm is bounded in $\mathcal{O}(nN)$
@@
The proof of the theorem is based on the well known bound on the prediction error of the LASSO algorithm of Theorem 5. Consider the algorithm described in Algorithm 1. First, we demonstrate now that Algorithm 1 satisfies the condition of Theorem 6. In our setting, the input space of the learned model $h$ is $\mathcal{X}=[-1,1]^{m}\subset\mathbb{R}^{m}$ as we consider $h$ a function of the $m-$dimensional feature vector $\phi(\bm{x})$. Clearly, $||\phi(\bm{x})||_{\infty}\leq 1$ for all $x\in\mathcal{X}$. The hypothesis class $\mathcal{C}$ of our algorithm is given by the set of function of the same form of the learned $h$, i.e. $\mathcal{C}=\{\phi(\bm{x})\in\mathcal{X}\rightarrow w\cdot\phi(\bm{x}):||w||_{1}\leq B\}$ with $B=poly(n)$. With respect to Theorem 6, we can also choose $M=poly(n)$ so that $|h(\bm{x}_{l})-y_{l}|<M$ for all $l=1,...,N$:
$|h(\bm{x}_{l})-y_{l}|\leq|w\cdot\phi(\bm{x}_{l})|+|y_{l}|\leq||w||_{1}||\phi(\bm{x}_{l})||_{\infty}+2\leq poly(n)$ (31)
@@
$\epsilon_{1}=\max_{\begin{subarray}{c}i\in[1,...,m]\\ \bm{x}\in\{0,1\}^{n}\end{subarray}}\;\;|[\rho(\bm{x})P_{i}]-\phi(\bm{x}_{i})|$ (33)
@@
$\leq(|\text{Tr}[\rho(\bm{x}_{\ell^{*}})O({\bm{\alpha}})]-h(\bm{x}_{\ell^{*}})|+|\text{Tr}[\rho(\bm{x}_{\ell^{*}})O({\bm{\alpha}})]-y_{\ell}|)^{2}$ (36)
@@
In practice, we can require to the LASSO algorithm to obtain a $w^{*}$ which achieves a training error at most $\epsilon_{3}/2$ larger than the optimal one. We can obtain such precision by setting $B=poly(n)$. Formally, we have
@@
Because we have set $B=poly(n)$, we have that the second term must be at most $\hat{R}(h_{opt})$ and therefore we have
@@
We note that as $B\leq\mathcal{O}(poly(n))$ we can bound the error $\epsilon^{\prime}_{1}=B\epsilon_{1}$ to scale polynomially to zero by just reducing the sampling error $\epsilon_{1}$ using polynomially many more copies of each $\rho(\bm{x}_{l})$. Thus we can rewrite equation (32) as
@@
$2BM\sqrt{\frac{2\log(2m)}{N}}+M^{2}\sqrt{\frac{\log(\delta)^{-1}}{2N}}\leq\left(\frac{B^{2}}{\sqrt{N}}(\sqrt{2\log(2\mathcal{O}(n))}+\sqrt{\log(\delta^{-1})}\right)$ (43)
@@
$N$ $=2\frac{B^{4}\sqrt{2\log(\mathcal{O}(n)/\delta)}}{\epsilon_{3}}$ (44)
$\sim\frac{\log(n/\delta)poly(n)}{\epsilon_{3}}$ (45)
By setting $\epsilon^{\prime}_{1}=0.2\epsilon$, $\epsilon_{2}=\epsilon$, and $\epsilon_{3}=0.4$ we have $(\epsilon^{\prime}_{1}+\epsilon_{2})^{2}+\epsilon_{3}\leq\epsilon$ and recover the claim of the Theorem.
Finally we bound the efficiency of our algorithm. The training time is dominated by the creation of the feature map $\phi(\bm{x}_{\ell})$ for each training point and by the LASSO regression over the corresponding feature space. To create the vector $\phi(\bm{x})$ the quantum algorithm needs to prepare multiple copies of $\rho(\bm{x}_{\ell})$ $\forall\bm{x}_{\ell}\in T_{N}^{\bm{\alpha}}$. As seen before, only a polynomial number of copies are sufficient to achieve a desired error $\epsilon^{\prime}_{1}$, so that the whole process takes time $\mathcal{O}(nN)$. For the LASSO regression, it is known that to obtain a training error at most $\epsilon_{3}/2$ larger than the optimal function value, the LASSO algorithm on the feature space of $\phi(\bm{x})$ can be executed in time $\mathcal{O}\left(\frac{m_{\phi}\log m_{\phi}}{\epsilon^{2}_{3}}\right)$ [HK12], where in our case $m_{\phi}=m$. It is easy to show that even this time is bounded by $\mathcal{O}(nN)$. The prediction time corresponds to the time for the evaluation of the learned model $h(\bm{x})=w^{*}\cdot\phi(\bm{x})$ which takes time $\mathcal{O}(m)\sim\mathcal{O}(n)$. In conclusion, the overall time of the quantum algorithm is bounded by $\mathcal{O}(nN)$
Appendix B Proof of Theorem 3 and Corollary 1

### v2

@@ A.2 Proof of Theorem 2
Let $\{U^{n}_{BQP}\}_{n}$ be a family of quantum circuits which decides an arbitrary $\operatorname{\mathsf{BQP}}$ language $\mathcal{L}$, one circuit per size. Precisely, according to the definition of $\operatorname{\mathsf{BQP}}$ in Def. 2, this implies that for any $\bm{x}\in\{0,1\}^{n}$ measuring the first qubit in the computational basis on the state $U^{n}_{BQP}\ket{\bm{x}}$ will output 1 (or 0) with probability greater that 2/3 if $\bm{x}\in\mathcal{L}$ ($\bm{x}\not\in\mathcal{L}$). It is obvious that the the quantum model $f^{O^{\prime}}=\text{Tr}[\rho(\bm{x})O^{\prime}]$ correctly decides every $\bm{x}\in\mathcal{L}$ if $\rho(\bm{x})=U^{n}_{BQP}\ket{\bm{x}}\bra{\bm{x}}(U^{n}_{BQP})^{\dagger}$ and $O^{\prime}=Z\otimes\underbrace{I\otimes...\otimes I}_{n-1}$. In fact measuring the first qubit of $U^{n}_{BQP}\ket{\bm{x}}$:
@@
Therefore, as $f^{O^{\prime}}(\bm{x})>0$ if $\bm{x}\in\mathcal{L}$ and $f^{O^{\prime}}(\bm{x})<0$ if $\bm{x}\notin\mathcal{L}$ such quantum model could efficiently decide the language $\mathcal{L}$. By Lemma 1, for every $n$ there exist a 4-local Hamiltonian $H_{\mathrm{\mathrm{hard}}}$ such that evolving the state |$\ket{\bm{x}}$ for a time $t=\pi$ under will produce the state $U^{n}_{\operatorname{\mathsf{BQP}}}\ket{\bm{x}}$ (on the work register). Therefore, the concept class $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$ contains a concept $f^{\bm{\alpha^{\prime}}}$, where $\bm{\alpha^{\prime}}$ is defined such that $O(\bm{\alpha^{\prime}})=O^{\prime}$, which correctly decides $\mathcal{L}$. The final step of the proof lies in noticing that by Lemma 2, if $\mathcal{F}_{\mathrm{evolved}}^{H_{\mathrm{hard}},O}$ is learnable in the sense of Def. 4, then there must exists a polynomial size classical circuit which evaluates $f^{\bm{\alpha^{\prime}}}$ correctly on every $\bm{x}\in\{0,1\}^{n}$. As $f^{\bm{\alpha}}$ decides the $\operatorname{\mathsf{BQP}}$ language $\mathcal{L}$, and for any $\operatorname{\mathsf{BQP}}$ language we can construct such Hamiltonian $H_{\mathrm{hard}}$, this implies that $\operatorname{\mathsf{BQP}}\subseteq\mathsf{P/poly}$.
Appendix B Proof of quantum learnability
@@
In this Section we prove that Algorithm 1 satisfies the learning condition in Def. 4 using only polynomial resources in both time and samples. As a first step, notice that the quanutm states $\rho_{H_{\mathrm{hard}}}(\bm{x})$ in Eq. (3) of $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$ are easily preparable on a quantum computer. This is due to the fact that the Hamiltonians $H_{\mathrm{hard}}$ used to establish our hardness result in Theorem 2 are local, allowing their time-evolution to be efficiently simulated on a quantum computer. Since the quantum algorithm can efficiently prepare time evolved states of local Hamiltonians, the quantum states $\rho_{H_{\mathrm{hard}}}(\bm{x})=U\ket{\bm{x}}\bra{\bm{x}}U^{\dagger}$ with $U=e^{iH_{\mathrm{\mathrm{hard}}}\pi}$ can efficiently be prepared on a quantum computer. In order to demonstrate the quantum learnability of $\mathcal{F}^{H_{\mathrm{hard}},O}_{\mathrm{evolved}}$, all that remains is to rigorously bound the sample and time complexity of Algorithm 1 in the following Theorem 7. As the bound obtained on the sample complexity derives from the generalization bound of the LASSO algorithm [17, 10], we first repeat this result.
Theorem 6 (Theorem 11.16 in [17]).
@@
Let $\mathcal{X}\subseteq\mathbb{R}^{A}$ and $\mathcal{C}=\{\bm{x}\in\mathcal{X}\mapsto\vec{w}\cdot\bm{x}:\|\vec{w}\|_{1}\leq B\}$. Let $\mathcal{S}=((\bm{x}_{1},y_{1}),\dots,(\bm{x}_{N},y_{N}))\in(\mathcal{X}\times\mathcal{Y})^{N}$. Let $\mathcal{D}$ denote a distribution over $\mathcal{X}\times\mathcal{Y}$ according to which the training data $\mathcal{S}$ is drawn. Assume that there exists $r_{\infty}>0$ such that for all $\bm{x}\in\mathcal{X}$, $\|\bm{x}\|_{\infty}\leq r_{\infty}$ and $M>0$ such that $|h(\bm{x})-y|\leq M$ for all $(\bm{x},y)\in\mathcal{X}\times\mathcal{Y}$. Then, for any $\delta>0$, with probability at least $1-\delta$, each of the following inequalities holds for all $h\in\mathcal{C}$:
$\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}[|h(\bm{x})-y|^{2}]:=R(h)$ (20)
$\leq\hat{R}_{\mathcal{S}}(h)+2r_{\infty}BM\sqrt{2\log(2|A|)N}+M\sqrt{2\log(1/\delta)/{2N}}$
@@
We are now ready to state Theorem 7 (a rigorous formulation of Lemma 3 in the main text) which provides precise guarantees on the number of samples and the time complexity required by Algorithm 1.
@@
Given $n$, $\delta>0$, $\frac{1}{e}>\epsilon>0$ and a training data set $\mathcal{T}^{\bm{\alpha}}_{N}=\{(\bm{x}_{\ell},y_{\ell})\}_{\ell=1}^{N}$ of size
$N=\mathcal{O}\left(\frac{\log(\mathrm{poly}(n)/\delta)\mathrm{poly}(n)}{\epsilon^{2}}\right)$ (21)
where $\bm{x}_{\ell}$ is sampled from an unknown distribution $\mathcal{D}$ over $\bm{x}\in\{0,1\}^{n}$ and $|y_{\ell}-\text{Tr}(O({\bm{\alpha}})\rho(\bm{x}_{\ell}))|\leq\epsilon$ for any geometrically local observable $O({\bm{\alpha}})=\sum_{i=1}^{m}{\bm{\alpha}}_{i}P_{i}$ with $P_{i}\in\{I,X,Y,Z\}^{\otimes n}$ and ${\bm{\alpha}}\in[-1,1]^{m}$, and $\rho(\bm{x})=U(\bm{x})\ket{0}\bra{0}U^{\dagger}(\bm{x})$. Then there exists a quantum algorithm $\mathcal{A}_{q}(\mathcal{T}^{\alpha}_{N},\bm{x})=h(\bm{x})$ such that:
@@
with probability at least $1-\delta$. The computational time of the quantum algorithm is bounded in $\mathcal{O}(\mathrm{poly}(n)N)$
@@
The proof of the theorem is based on the well known bound on the prediction error of the LASSO algorithm of Theorem 6. Consider the algorithm described in Algorithm 1. First, we demonstrate now that Algorithm 1 satisfies the condition of Theorem 6. In our setting, the input space of the learned model $h$ is $\mathcal{X}=[-1,1]^{m}\subset\mathbb{R}^{m}$ as we consider $h$ a function of the $m-$dimensional feature vector $\phi(\bm{x})$. Clearly, $||\phi(\bm{x})||_{\infty}\leq 1$ for all $\bm{x}\in\mathcal{X}$. The hypothesis class $\mathcal{C}$ of our algorithm is given by the set of functions of the same form of the learned $h$, i.e. $\mathcal{C}=\{\phi(\bm{x})\in\mathcal{X}\rightarrow w\cdot\phi(\bm{x}):||w||_{1}\leq B\}$ with $B=\mathrm{poly}(n)$. With respect to Theorem 6, we can also choose $M=\mathrm{poly}(n)$ so that $|h(\bm{x}_{l})-y_{l}|<M$ for all $l=1,...,N$:
$|h(\bm{x}_{l})-y_{l}|$ $\leq|w\cdot\phi(\bm{x}_{l})|+|y_{l}|\leq||w||_{1}||\phi(\bm{x}_{l})||_{\infty}+2$ (23)
$\leq\text{poly}(n)$
@@
$\epsilon_{1}=\max_{\begin{subarray}{c}i\in[1,...,m]\\ \bm{x}\in\{0,1\}^{n}\end{subarray}}\;\;|\mathrm{Tr}[\rho(\bm{x})P_{i}]-\phi(\bm{x}_{i})|$ (25)
@@
$\leq(|\text{Tr}[\rho(\bm{x}_{\ell^{*}})O({\bm{\alpha}})]-h(\bm{x}_{\ell^{*}})|\;+$ (28)
$\;\;\;\;\;\;|\text{Tr}[\rho(\bm{x}_{\ell^{*}})O({\bm{\alpha}})]-y_{\ell}|)^{2}$
@@
In practice, we can require to the LASSO algorithm to obtain a $w^{*}$ which achieves a training error at most $\epsilon_{3}/2$ larger than the optimal one. We can obtain such precision by setting $B=\mathrm{poly}(n)$. Formally, we have
@@
Because we have set $B=\mathrm{poly}(n)$, we have that the second term must be at most $\hat{R}(h_{opt})$ and therefore we have
@@
We note that as $B\leq\mathcal{O}(\mathrm{poly}(n))$ we can bound the error $\epsilon^{\prime}_{1}=B\epsilon_{1}$ to scale polynomially to zero by just reducing the sampling error $\epsilon_{1}$ using polynomially many more copies of each $\rho(\bm{x}_{l})$. Thus we can rewrite equation (24) as
@@
$2BM\sqrt{\frac{2\log(2m)}{N}}+M^{2}\sqrt{\frac{\log(\delta)^{-1}}{2N}}$
$\leq\left(\frac{B^{2}}{\sqrt{2N}}(4\sqrt{\log(2\mathcal{O}(\text{poly}(n)))}+\sqrt{\log(\delta^{-1})}\right)$ (35)
@@
$N$ $=2\frac{B^{4}\sqrt{2\log(\mathcal{O}(\text{poly}(n))/\delta)}}{(\epsilon_{3})^{2}}$ (36)
$\sim\frac{\log(\text{poly}(n)/\delta)\text{poly}(n)}{(\epsilon_{3})^{2}}$ (37)
By setting $\epsilon^{\prime}_{1}=0.2\epsilon$, $\epsilon_{2}=\epsilon$, and $\epsilon_{3}=0.4\epsilon$ we have $(\epsilon^{\prime}_{1}+\epsilon_{2})^{2}+\epsilon_{3}\leq\epsilon$ and recover the claim of the Theorem.
Finally we bound the efficiency of our algorithm. The training time is dominated by the creation of the feature map $\phi(\bm{x}_{\ell})$ for each training point and by the LASSO regression over the corresponding feature space. To create the vector $\phi(\bm{x})$ the quantum algorithm needs to prepare multiple copies of $\rho(\bm{x}_{\ell})$ $\forall\bm{x}_{\ell}\in T_{N}^{\bm{\alpha}}$. As seen before, only a polynomial number of copies are sufficient to achieve a desired error $\epsilon^{\prime}_{1}$, so that the whole process takes time $\mathcal{O}(\mathrm{poly}(n)N)$. For the LASSO regression, it is known that to obtain a training error at most $\epsilon_{3}/2$ larger than the optimal function value, the LASSO algorithm on the feature space of $\phi(\bm{x})$ can be executed in time $\mathcal{O}\left(\frac{m_{\phi}\log m_{\phi}}{\epsilon^{2}_{3}}\right)$ [54], where in our case $m_{\phi}=m$. It is easy to show that even this time is bounded by $\mathcal{O}(\mathrm{poly}(n)N)$. The prediction time corresponds to the time for the evaluation of the learned model $h(\bm{x})=w^{*}\cdot\phi(\bm{x})$ which takes time $\mathcal{O}(m)\sim\mathcal{O}(\mathrm{poly}(n))$. In conclusion, the overall time of the quantum algorithm is bounded by $\mathcal{O}(\mathrm{poly}(n)N)$
Appendix C Construction of the Kitaev’s circuit-to-Hamiltonian for the ground state
To make our work more self-contained, we report the construction of the Kitaev’s circuit-to-Hamiltonian which was used in the proof of Theorem 4. We follow the constructions outlined in [29, 30]. Let $x\in\mathcal{L}$ an input of a language $\mathcal{L}\in\operatorname{\mathsf{BQP}}$ and $U(x)=U_{T}...U_{1}$ be a quantum circuit of polynomial size $T$ acting on $n=\mathrm{poly}(|x|)$ qubits which decides $x$. We first add $2T$ layers of identities to $U(x)$. The Hamiltonian $H$ that is constructed operates on a space of $n=N+\log(3T+1)$ qubits. The first $N$ qubits represent the computation, and the last $\log(3T+1)$ qubits represent the possible values $0,\dots,3T$ for the clock. The Hamiltonian is constructed from three terms:
$H(x)=H_{\mathrm{init}}+H_{\mathrm{clock}}+\sum_{t=1}^{3T}H_{t}(x),$ (38)
with
$\centering H_{\mathrm{init}}\@add@centering$ $=\sum_{i=1}^{N}\ket{0}\bra{0}_{i},\quad H_{\mathrm{clock}}=\sum_{t=1}^{3T-1}\ket{01}\bra{01}^{\mathrm{clock}}_{t,t+1},$ (39)
$H_{t}(x)=\frac{1}{2}\Big{(}$ $I\otimes\ket{100}\bra{100}^{\mathrm{clock}}_{t-1,t,t+1}$ (40)
$+I\otimes\ket{110}\bra{110}^{\mathrm{clock}}_{t-1,t,t+1}$ (41)
$-U_{t}(x)\otimes\ket{110}\bra{100}^{\mathrm{clock}}_{t-1,t,t+1}$ (42)
$-U_{t}(x)^{\dagger}\otimes\ket{100}\bra{110}^{\mathrm{clock}}_{t-1,t,t+1}\Big{)}$ (43)
where |..|$\ket{.}\bra{.}_{i}$ acts on the $i$th site of $\mathbb{C}^{2^{N}}$, |..|$\ket{.}\bra{.}_{j}^{\mathrm{clock}}$ acts on the $j$th site of $\mathbb{C}^{2^{3T}}$ and $U_{t}$ denotes the $t$th layer of gates in $U(x)$. Note that $H(x)$ is 5-local for all $x$. The ground state of $H(x)$ is given by $\rho(x)=\ket{\psi(x)}\bra{\psi(x)}$ |, where
|$\ket{\psi(x)}=\frac{1}{\sqrt{3T}}\sum_{t=1}^{3T}(U_{t}\cdots U_{1})(x)\ket{0^{N}}\ket{1^{t}0^{3T-t}},$ (44)
Measuring the local observable $O=Z\otimes I\otimes\ket{1}\bra{1}_{T}^{\mathrm{clock}}$ on the state |$\ket{\psi(x)}$ will decide the input $x$, analogously to the case in the proof of Theorem 2.
Appendix D Proof of Theorem 5 and Corollary 1